OpenStudy (anonymous):

Use the function defined below to find an equation of the tangent line at the point indicated. y = cos(x), x = pi/3

OpenStudy (anonymous):

Please show steps.

OpenStudy (anonymous):

take the derivative, get \[-\sin(x)\] then find \[-\sin(\frac{\pi}{3})\] for your slope.

OpenStudy (anonymous):

then use the point - slope formula to find the equation for the line.

OpenStudy (anonymous):

Would I write it as -sin(pi/3) or would I make it -(sqrt(3)/2)?

OpenStudy (anonymous):

you need the second one, the actual number

OpenStudy (anonymous):

Alright. Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Could you help me finish the equation? :p